# iTACO: Interactable Digital Twins of Articulated Objects from Casually Captured RGBD Videos

### v2

@@ 1 Introduction
Building digital twins of articulated objects has many applications in computer animation, embodied AI, and robotics. For example, scalable pipelines for building interactable digital twins benefit robotics research, by enabling convenient training and evaluation in simulation prior to real-world deployment. However, current approaches to building interactable replicas of articulated objects impose many constraints on the input data, preventing practical and efficient acquisition of articulated objects in realistic settings. Thus, researchers still largely rely on manual annotation of articulation joint parameters [53].
Recent approaches to articulated object reconstruction and motion analysis can be categorized into three families. The first focuses on dynamic reconstruction of the shape of articulated objects [34, 78, 54, 62, 44]. Unfortunately, the final output is not interactable as it is typically a sequence of meshes instead of a kinematic structure with explicit links and joints. The second family infers joint parameters and builds part-level geometry simultaneously from two different states of the object [12, 24, 64, 27, 29, 36]. However, this requires aligning observations between states, and even small errors in alignment can lead to dramatic drops in accuracy. The third family of approaches generates articulated objects leveraging large language and multimodal models [30, 18]. At present, these approaches either rely on an external library for object retrieval or require data for fine-tuning. Retrieving objects is not easily scalable, while collecting data for fine-tuning is labor-intensive. Furthermore, generalization to real settings after fine-tuning on limited data is not easy.
By analyzing the issues of recent approaches, we summarize three desirable features for practical interactable digital twin acquisition: 1) minimal constraints on input; 2) explicit joint parameter and part-level geometry representation for easy integration into simulators; and 3) generalization across diverse object categories. To address these desiderata, we create digital twins of articulated objects from casually captured RGBD videos showing a human interacting with the object. For example, a human operator with a hand-held camera can first scan the object to reconstruct the object surface, and then record an interaction with the object. Such videos are easy to capture at scale and contain the necessary motion information to infer joint parameters. Compared to the problem setting in prior work, casually captured videos are far less constrained and easier to capture, the final output is easier to use for simulation, and the overall pipeline is scalable to real-world settings.
@@
Since our problem setting has not been previously addressed, we build a dataset for evaluation. We generate videos for 284 synthetic objects from 11 categories in PartNet-Mobility [33] and collect real casual RGBD videos. Our dataset contains 20 more objects compared to prior works [24, 27], and allows a more robust evaluation of articulated object acquisition. In summary, we make the following contributions:
@@ 2 Related Work
Our work is related to motion analysis and reconstruction of articulated objects, and dynamic scene understanding. We review recent work in both fields.
@@ 2.1 Articulated Object Reconstruction
Motion analysis and reconstruction of articulated objects is a long-standing task in computer vision and robotics. Previous motion analysis works focus on predicting joint parameters or part-level segmentation from an image or a single point cloud [60, 74, 20, 63, 11, 48, 81]. These works require training on specific datasets, which limits generalization. Previous articulated object reconstruction works focus on reconstructing the shape of articulated objects [34, 77, 78, 62, 44]. The final output of these methods is the entire shape of the object, which is not interactable, in particular due to the absence of a kinematic structure with joints and articulation parameters that can be driven kinematically or through physics-based simulation. In contrast, we aim to build interactable digital twins of articulated objects, where inferring joint parameters and part-level geometry is essential.
@@
Another line of work attempts to infer joint parameters and build part-level geometry of articulated objects simultaneously. One problem setting uses multi-view observations of different object states [12, 24, 64, 27]. These works implicitly assume all observations have been aligned to the same coordinates. In practice, this requires either manual alignment which is not scalable or algorithmic alignment which is error-prone. Other prior work uses robots to interact with the articulated object and reconstruct articulated objects from the interaction [29, 36, 75, 61]. However, robots are expensive and may damage themselves or the environment during the interaction. Therefore, this is not a practically scalable path to build digital twins of articulated objects. Recently, researchers began exploring motion analysis from videos [14, 65, 85]. Robot-See-Robot-Do [14] recovers 3D motions of an articulated object from a monocular human demonstration video. Their problem setting is close to us, but requires the camera to be static.
@@
Other work uses large language and multimodal models to generate a code representation or kinematic graph from a single image of an articulated object [30, 18, 4, 25]. Leveraging vision-language models, Articulate-Anything can also take image or video as input [18]. These works either rely on existing mesh libraries to retrieve object part meshes or require extra data for fine-tuning foundation models. Thus, at the current stage, this line of work still suffers from out-of-distribution issues, limiting generalization and scalability. In contrast, our method does not rely on object retrieval or fine-tuning, and is inherently more general and scalable.
@@
In computer animation, automatic skeletal rigging has been studied for decades [49, 1, 72, 55, 73, 70, 19, 71, 47, 45, 5, 23, 83]. However, the skeleton model is designed to animate humans and animals, whose links are non-rigid and joints are mainly ball joints. For rigid articulated objects such as drawers and microwaves, most joints are prismatic joints and revolute joints. Thus, it is suitable to represent rigid articulated objects using skeleton models.
@@ 2.2 Dynamic Scene Reconstruction
Dynamic scene reconstruction is a challenging task in computer vision. One line of research describes a dynamic scene with a deformation field. Signed distance functions, and later neural radiance fields and 3D gaussian splatting are commonly used [35, 2, 43, 38, 6, 80, 28, 66, 56, 69, 59]. These works achieve impressive results in shape reconstruction and novel view rendering of dynamic scenes, but it is difficult to transform the deformation field to an explicit motion representation, such as a rigid transformation matrix, which is necessary for our problem.
Another line of approaches attempts to reconstruct dynamic scenes from monocular depth estimation and optical flow. Robust-CVD proposes to first estimate the depth of each frame and then optimize camera poses and depth jointly [17]. CasualSAM leverages optical flow to train a depth and movement prediction network and optimize camera parameters [84]. MegaSaM proposes an optimization framework for dynamic scene reconstruction based on DROID-SLAM [21, 52]. Recently, DUSt3R [58] provides a powerful foundation model for static scene reconstruction. There are several recent extensions to DUSt3R. MonST3R fine-tunes DUSt3R on dynamic scene datasets to recover dynamic 3D scenes [82]. CUT3R introduces a continuous 3D perception model for online dense 3D reconstruction [57]. Easi3R identifies moving objects in the video from DUSt3R’s attention maps [3]. Aether proposes a unified model for reconstruction, planning, and prediction in dynamic scenes [50]. Our coarse prediction pipeline directly benefits from MonST3R [82], and our refinement framework is inspired by the optimization framework in CasualSAM [84] and Mega-SaM [21].
A key sub-task in dynamic scene understanding is segmenting the ‘moving map’ (mask image indicating which pixels are moving in the scene at each frame in the video). Recent approaches are mainly based on optical flow or point tracking [76, 32, 31, 68, 13, 8, 10]. However, we find that point tracking methods struggle with textureless objects and are sensitive to point locations (see supplement). Therefore, we do not use these approaches.
@@ 3.2 Coarse Prediction
In practice, we use MonST3R [82] to provide a rough moving map of each video frame. Then we unproject pixels in to 3D point cloud with camera intrinsics, and use LoFTr [46] to compute pixel-level feature matching between two video frames. We use feature matching in static regions identified by to compute relative camera poses. We align all observations to the same camera coordinate system with the relative camera poses of each frame. Then, we use feature matching in the dynamic regions to compute a rigid transformation of the movable parts of the object. We can further estimate joint parameters from this rigid transformation. We estimate joint parameters from multiple video frame pairs and average the results as the coarse prediction of the joint parameters and states.
@@ 3.3 Refinement
To tackle the second problem our automatic part segmentation module aggregates pixels into object parts. This is more efficient and effective as the number of parts is much smaller than the number of pixels, reducing optimization complexity. Similar to Segment-Anything [16], automatic part segmentation segments the first video frame into several part regions, and then tracks these regions throughout the video using SAM2 [39]. Subsequently, instead of optimizing on each pixel, we optimize the probability of being a moving part for each tracked segment. We define a vector to represent the probability of each segment being a moving part, where is the number of segments. To construct for a frame, we multiply each probability value by the corresponding mask of the part. Finally, we sum the results across different parts to construct the moving map. We leverage an existing implementation of object part segmentation for videos.
@@
Since our pipeline only estimates the moving map of the input video, we need to find the movable part of the object on . We unproject the pixels in the moving region in the moving map of the first frame . Then, we select points on which are close to the unprotected moving pixels to be the movable part of the object. We use NKSR [9] to build meshes of both the static and the movable parts of the object.
@@
| Revolute | Articulate Anything | 0.820.79 | 0.810.40 | 40.90 | N.A. | 38.63 |
| RSRD | 1.160.52 | 1.181.21 | 88.88 | 1.050.59 | 46.66 | |
@@
| Prismatic | Articulate Anything | 0.920.78 | N.A. | 50.00 | N.A. | 46.66 |
| RSRD | 1.270.44 | N.A. | 40.00 | 0.630.41 | 36.66 | |
@@ 4.1 Experiment Setup
At present, there is no benchmark for evaluating joint parameter estimation from dynamic RGBD videos. Therefore, we construct a synthetic dataset for evaluation. We select objects from 11 categories in the PartNet-Mobility dataset [33]. We simulate interaction between human hands and these objects using the SAPIEN simulator [67]. For each interaction, we generate two videos shot from two different viewpoints. Camera trajectories are generated by moving the camera towards a new target pose every 7 - 10 timesteps in the simulator. The target pose is sampled from a Gaussian distribution, where the mean value is the current camera pose and the standard deviation is set to ensure the camera can consistently look at the object. To get the object surface point cloud, we unproject the depth maps from 24 viewpoints to 3D and fuse the point cloud. For Robot-See-Robot-Do, we generate 100 views because it requires producing a NeRF and 3D Gaussian splat [14]. Please refer to the supplement for more details about the dataset construction. Since Robot-See-Robot-Do requires manually cropping and clustering 3D gaussian splats which is labor-intensive, we split the dataset into two parts. We randomly sampled 10% of test videos from our dataset to construct a smaller dataset named S-Dataset, containing 73 test videos, while L-Dataset contains the remaining videos in the original dataset.
@@
Robot-See-Robot-Do (RSRD) [14]: a differentiable rendering pipeline that recovers 3D part motion from human demonstration videos. We follow the pipeline to first reconstruct and segment the 3D Gaussian splats of the object. Then, the input video is used to estimate the SE3 transformation of the 3D Gaussians for each part at each frame via differentiable rendering. For a fair comparison, we modify their tracking pipeline to exploit ground truth depth maps of the input video. We select the group of 3D Gaussians with the longest moving distance to be the moving part of the object. Since this pipeline does not include a geometry reconstruction module, we export 3D Gaussians and reconstruct mesh using NKSR [9] for geometry reconstruction.
Articulate Anything [18]: a system that retrieves object meshes from a library and estimates link and joint parameters from text, image, or video input. We follow the original work and use the Partnet-Mobility dataset as the object mesh library, which contains all the ground truth object meshes for our test cases, and use Gemini Flash-1.5 [51] for the VLM module.
@@
We use a suite of metrics to evaluate both geometric reconstruction and articulation parameter estimation. Geometric accuracy is evaluated using Chamfer Distance between ground truth and estimated geometry for the whole object (CD-w), for the movable part (CD-m) and for the static part (CD-s). Articulation parameter estimation is evaluated using joint type classification error (percentage), joint axis error (radians), joint position error for revolute joints (meters), and joint state error (radians for revolute joints and meters for prismatic joints). If the method crashes or retrieves incorrect objects (for methods based on retrieval), we classify the output as a failure and assign error for the joint axis and state of revolute joints, and error equal to 1 for the other metrics.
@@ 4.2 Articulation Parameter Estimation Results
| Articulate Anything | 0.110.22 | 0.590.73 | 0.070.18 |
| RSRD | 3.3921.50 | 0.600.60 | 0.170.44 |
@@
The quantitative results on S-Dataset are in Tab.˜1, and qualitative results are in Fig.˜4. Results on L-Dataset can be found in the supplement. Our iTACO method outpeforms or is comparable to state-of-the-art methods along most metrics. Articulate Anything struggles with retrieving the correct object meshes. Most failure cases in Tab.˜1 are due to retrieval errors, which is surprising since the retrieval library contains all the ground truth object meshes. For example, in Fig.˜4 we find that Articulate Anything retrieves a lamp in the Stapler test case. In addition, the link placement and affordance detection modules appear to not be robust either, showing the challenge of our problem setting. Articulate Anything selects the wheel of the table to be a movable part, and places the lid of the USB stick below the main body.

### v3

@@ 1 Introduction
Building digital twins of articulated objects has many applications in computer animation, embodied AI, and robotics. For example, scalable pipelines for building interactable digital twins benefit robotics research, by enabling convenient training and evaluation in simulation prior to real-world deployment. However, current approaches to building interactable replicas of articulated objects impose many constraints on the input data, preventing practical and efficient acquisition of articulated objects in realistic settings. Thus, researchers still largely rely on manual annotation of articulation joint parameters [torne2024rialto].
Recent approaches to articulated object reconstruction and motion analysis can be categorized into three families. The first focuses on dynamic reconstruction of the shape of articulated objects [mu2021learning, yang2022banmo, tseng2022cla, wei2022self, song2024reacto]. Unfortunately, the final output is not interactable as it is typically a sequence of meshes instead of a kinematic structure with explicit links and joints. The second family infers joint parameters and builds part-level geometry simultaneously from two different states of the object [jiang2022ditto, jiayi2023paris, weng2024neural, liu2025building, lv2022sagci, nie2023structure]. However, this requires aligning observations between states, and even small errors in alignment can lead to dramatic drops in accuracy. The third family of approaches generates articulated objects leveraging large language and multimodal models [mandi2024real2code, le2024articulate]. At present, these approaches either rely on an external library for object retrieval or require data for fine-tuning. Retrieving objects is not easily scalable, while collecting data for fine-tuning is labor-intensive. Furthermore, generalization to real settings after fine-tuning on limited data is not easy.
By analyzing the issues of recent approaches, we summarize three desirable features for practical interactable digital twin acquisition: 1) minimal input constraints; 2) explicit representation of joint parameter and part-level geometry for easy integration into simulators; and 3) generalization across diverse object categories. To address these desiderata, we create digital twins of articulated objects from casually captured RGBD videos showing a human interacting with the object. For example, a human operator with a hand-held camera can first scan the object to reconstruct the object surface, and then record an interaction with the object. Such videos are easy to capture at scale and contain the necessary motion information to infer joint parameters. Compared to the problem setting in prior work, casually captured videos are far less constrained and easier to capture, the final output is easier to use for simulation, and the overall pipeline is scalable to real-world settings.
@@
Since our problem has not been previously addressed, we build a dataset for evaluation. We generate videos for 284 synthetic objects from 11 categories in PartNet-Mobility [Mo_2019_CVPR] and collect real RGBD videos. Our dataset contains 20 more objects compared to prior works [jiayi2023paris, liu2025building], allowing a more robust evaluation of articulated object acquisition. In summary, we make the following contributions:
@@ 2 Related Work
Our work is related to motion analysis and reconstruction of articulated objects, and dynamic scene understanding.
@@ 2.1 Articulated Object Reconstruction
Motion analysis and reconstruction of articulated objects is a long-standing task in computer vision and robotics. Previous motion analysis works focus on predicting joint parameters or part-level segmentation from an image or a single point cloud [wang2019shape2motion, Yan2019RPMNet, li2019category, weng2021captra, jiang2022opd, sun2023opdmulti, yu2024gamma]. These works require training on specific datasets, which limits generalization. Previous articulated object reconstruction works focus on reconstructing the shape of articulated objects [mu2021learning, yang2021lasr, yang2022banmo, wei2022self, song2024reacto]. The final output of these methods is the entire shape of the object, which is not interactable, in particular due to the absence of a kinematic structure with joints and articulation parameters that can be driven kinematically or through physics-based simulation. In contrast, we aim to build interactable digital twins of articulated objects, where inferring joint parameters and part-level geometry is essential.
@@
Another line of work attempts to infer joint parameters and build part-level geometry of articulated objects simultaneously. One problem setting uses multi-view observations of different object states [jiang2022ditto, jiayi2023paris, weng2024neural, liu2025building]. These works implicitly assume all observations have been aligned to the same coordinates. In practice, this requires either manual alignment which is not scalable or algorithmic alignment which is error-prone. Other prior work uses robots to interact with the articulated object and reconstruct articulated objects from the interaction [lv2022sagci, nie2023structure, InterRecon23, wang2024articulated]. However, robots are expensive and may damage themselves or the environment during the interaction. Therefore, this is not a practically scalable path to build digital twins of articulated objects. Recently, researchers began exploring motion analysis from videos [kerr2024rsrd, werby2025articulated, zhou2025monomobility]. Robot-See-Robot-Do [kerr2024rsrd] recovers 3D motions of an articulated object from a monocular human demonstration video. Their problem setting is close to us, but requires the camera to be static.
@@
Other work uses large language and multimodal models to generate a code representation or kinematic graph from a single image of an articulated object [mandi2024real2code, le2024articulate, chen2024urdformer, jiayi2024singapo]. Leveraging vision-language models, Articulate-Anything can also take image or video as input [le2024articulate]. These works either rely on existing mesh libraries to retrieve object part meshes or require extra data for fine-tuning foundation models. Thus, at the current stage, this line of work still suffers from out-of-distribution issues, limiting generalization and scalability. In contrast, our method does not rely on object retrieval or fine-tuning, and is inherently more general and scalable.
@@
In computer animation, automatic skeletal rigging has been studied for decades [Tagliasacchi2009curve, Baran2007auto, yan2016erosion, Tzionas2016reconstructing, yan2018voxel, xu2020rignet, li2021nbs, xu2022morig, sun2025drive, Song2025magicarticulate, deng2025anymate, liu2025riganything, zhang2025unirig]. However, the skeleton model is designed to animate humans and animals, whose links are non-rigid and joints are mainly ball joints. For rigid articulated objects such as drawers and microwaves, most joints are prismatic joints and revolute joints. Thus, it is suitable to represent rigid articulated objects using skeleton models.
@@ 2.2 Dynamic Scene Reconstruction
Dynamic scene reconstruction is a challenging task in computer vision. One line of research describes a dynamic scene with a deformation field. Signed distance functions, and later neural radiance fields and 3D gaussian splatting are commonly used [newcombe2015dynamicfusion, bozic2020deepdeform, slavcheva2018sobolevfusion, pumarola2020d, duan:2024:4drotorgs, yang2023gs4d, luiten2023dynamic, Wu_2024_CVPR, som2024, xu2024das3r, wang2025gflow]. These works achieve impressive results in shape reconstruction and novel view rendering of dynamic scenes, but it is difficult to transform the deformation field to an explicit motion representation, such as a rigid transformation matrix, which is necessary for our problem.
Another line of approaches attempts to reconstruct dynamic scenes from monocular depth estimation and optical flow. Robust-CVD proposes to first estimate the depth of each frame and then optimize camera poses and depth jointly [kopf2021rcvd]. CasualSAM leverages optical flow to train a depth and movement prediction network and optimize camera parameters [CasualSaM]. MegaSaM proposes an optimization framework for dynamic scene reconstruction based on DROID-SLAM [li2024_MegaSaM, teed2021droid]. Recently, DUSt3R [dust3r_cvpr24] provides a powerful foundation model for static scene reconstruction. There are several recent extensions to DUSt3R. MonST3R fine-tunes DUSt3R on dynamic scene datasets to recover dynamic 3D scenes [zhang2024monst3r]. CUT3R introduces a continuous 3D perception model for online dense 3D reconstruction [cut3r]. Easi3R identifies moving objects in the video from DUSt3R’s attention maps [chen2025easi3r]. Aether proposes a unified model for reconstruction, planning, and prediction in dynamic scenes [aether]. Our coarse prediction pipeline directly benefits from MonST3R [zhang2024monst3r], and our refinement framework is inspired by the optimization framework in CasualSAM [CasualSaM] and MegaSaM [li2024_MegaSaM].
A key sub-task in dynamic scene understanding is segmenting the ‘moving map’ (mask image indicating which pixels are moving in the scene at each frame in the video). Recent approaches are mainly based on optical flow or point tracking [yang2021selfsupervised, meunier2023em, meunier2023unsupervised, xie2024flowsam, karazija24learning, golisabour2024romo, huang2025segmentmotionvideos]. However, we find that point tracking methods struggle with textureless objects and are sensitive to point locations (see supplement). Therefore, we do not use these approaches.
@@ 3.2 Coarse Prediction
In practice, we use MonST3R [zhang2024monst3r] to provide a rough moving map of each video frame. Then we unproject pixels in to 3D point cloud with camera intrinsics, and use LoFTr [sun2021loftr] to compute pixel-level feature matching between two video frames. We use feature matching in static regions identified by to compute relative camera poses. We align all observations to the same camera coordinate system with the relative camera poses of each frame. Then, we use feature matching in the dynamic regions to compute a rigid transformation of the movable parts of the object. We can further estimate joint parameters from this rigid transformation. We estimate joint parameters from multiple video frame pairs and average the results as the coarse prediction of the joint parameters and states.
@@ 3.3 Refinement
To tackle the second problem our automatic part segmentation module aggregates pixels into object parts. This is more efficient and effective as the number of parts is much smaller than the number of pixels, reducing optimization complexity. Similar to Segment-Anything [kirillov2023segment], automatic part segmentation segments the first video frame into several part regions, and then tracks these regions throughout the video using SAM2 [ravi2024sam2]. Subsequently, instead of optimizing on each pixel, we optimize the probability of being a moving part for each tracked segment. We define a vector to represent the probability of each segment being a moving part, where is the number of segments. To construct for a frame, we multiply each probability value by the corresponding mask of the part. Finally, we sum the results across different parts to construct the moving map. We leverage an existing implementation of object part segmentation for videos.
@@
Since our pipeline only estimates the moving map of the input video, we need to find the movable part of the object on . We unproject the pixels in the moving region in the moving map of the first frame . Then, we select points on which are close to the unprotected moving pixels to be the movable part of the object. We use NKSR [huang2023nksr] to build meshes of both the static and the movable parts of the object.
@@
| Revolute | AA [le2024articulate] | 0.820.79 | 0.810.40 | 40.90 | N.A. | 38.63 |
| RSRD [kerr2024rsrd] | 1.160.52 | 1.181.21 | 88.88 | 1.050.59 | 46.66 | |
@@
| Prismatic | AA [le2024articulate] | 0.920.78 | N.A. | 50.00 | N.A. | 46.66 |
| RSRD [kerr2024rsrd] | 1.270.44 | N.A. | 40.00 | 0.630.41 | 36.66 | |
@@ 4.1 Experiment Setup
At present, there is no benchmark for evaluating joint parameter estimation from dynamic RGBD videos. Therefore, we construct a synthetic dataset for evaluation. We select objects from 11 categories in the PartNet-Mobility dataset [Mo_2019_CVPR]. We simulate interaction between human hands and these objects using the SAPIEN simulator [Xiang_2020_SAPIEN]. For each interaction, we generate two videos shot from two different viewpoints. Camera trajectories are generated by moving the camera towards a new target pose every 7 - 10 timesteps in the simulator. The target pose is sampled from a Gaussian distribution, where the mean value is the current camera pose and the standard deviation is set to ensure the camera can consistently look at the object. To get the object surface point cloud, we unproject the depth maps from 24 viewpoints to 3D and fuse the point cloud. For Robot-See-Robot-Do, we generate 100 views because it requires producing a NeRF and 3D Gaussian splat [kerr2024rsrd]. Please refer to the supplement for more details about the dataset construction. Since Robot-See-Robot-Do requires manually cropping and clustering 3D gaussian splats which is labor-intensive, we split the dataset into two parts. We randomly sampled 10% of test videos from our dataset to construct a smaller dataset named S-Dataset, containing 73 test videos, while L-Dataset contains the remaining videos in the original dataset.
@@
Robot-See-Robot-Do (RSRD) [kerr2024rsrd]: a differentiable rendering pipeline that recovers 3D part motion from human demonstration videos. We follow the pipeline to first reconstruct and segment the 3D Gaussian splats of the object. Then, the input video is used to estimate the SE3 transformation of the 3D Gaussians for each part at each frame via differentiable rendering. For a fair comparison, we modify their tracking pipeline to exploit ground truth depth maps of the input video. We select the group of 3D Gaussians with the longest moving distance to be the moving part of the object. Since this pipeline does not include a geometry reconstruction module, we export 3D Gaussians and reconstruct a mesh using NKSR [huang2023nksr] for geometry reconstruction.
Articulate Anything (AA) [le2024articulate]: a system that retrieves object meshes from a library and estimates link and joint parameters from text, image, or video input. We follow the original work and use the Partnet-Mobility dataset as the object mesh library, which contains all the ground truth object meshes for our test cases, and use Gemini Flash-1.5 [team2023gemini] for the VLM module.
@@
We use a suite of metrics to evaluate both geometric reconstruction and articulation parameter estimation. Geometric accuracy is evaluated using Chamfer Distance between ground truth and estimated geometry for the whole object (CD-w), for the movable part (CD-m) and for the static part (CD-s). Articulation parameter estimation is evaluated using joint type classification error (percentage), joint axis error (radians), joint position error for revolute joints (meters), and joint state error (radians for revolute joints and meters for prismatic joints). If the method crashes or retrieves incorrect objects (for retrieval methods), we classify the output as a failure and assign error for joint axis and revolute joint state, and error equal to 1 for other metrics.
@@ 4.2 Articulation Parameter Estimation Results
| AA [le2024articulate] | 0.110.22 | 0.590.73 | 0.070.18 |
| RSRD [kerr2024rsrd] | 3.3921.50 | 0.600.60 | 0.170.44 |
@@
The quantitative results on S-Dataset are in Tab.˜1, and qualitative results are in Fig.˜4. Results on L-Dataset are in the supplement. Our iTACO method outperforms or is comparable to state-of-the-art methods along most metrics. AA struggles with retrieving the correct object meshes. Most failure cases are due to retrieval errors, which is surprising since the retrieval library contains all the ground truth object meshes. For example, in Fig.˜4 AA retrieves a lamp in the Stapler test case. In addition, the link placement and affordance detection modules appear to not be robust either, showing the challenge of our problem setting. AA selects the wheel of the table to be a movable part, and places the lid of the USB stick below the main body.